OpenStudy (anonymous):

For questions 1-2, what are the excluded values? 1. y = 4/9x-45 A.-5 B.0.2 C. 5 D. -0.2 2. y= 7/2x + 48 A.0.04 B.24 C.-24 D. -0.04

hartnn (hartnn):

Since, zero in the denominator makes the equation undefined, the values which make the denominator =0 will be excluded. so, for 1)st what vlue of x make 9x-45 =0 ?

OpenStudy (anonymous):

C @hartnn

hartnn (hartnn):

thats correct! :) in same way, for 2)nd 2x-48=0 for which value of x ?

OpenStudy (anonymous):

B @hartnn

hartnn (hartnn):

oh sorry, its 2x+48 so 2x+48 = 0 gives x = -24 , option C :)